OpenStudy (anonymous):

9. Find all the zeroes of the equation. x3 – x2 = –3x – 3 (1 point)

OpenStudy (anonymous):

\[x ^{3}-x ^{2}=-3x-3?????????\]

OpenStudy (anonymous):

Start by setting the equation equal to 0 x3 – x2 = –3x – 3 move –3x – 3 to the left side so: x3 – x2+3x-3=0

OpenStudy (anonymous):

So then I set those numbers equal to zero?

OpenStudy (anonymous):

Now, From x3 – x2+3x-3=0, you want to factor So (x-1)(x^2+3)=0 (x-1)=0 (x^2+3)=0

OpenStudy (campbell_st):

slight error the equation when set to zero is x^3 - x^2 + 3x + 3 = 0

OpenStudy (anonymous):

oops. My bad. So then just factor: x^3 - x^2 + 3x + 3 = 0

OpenStudy (campbell_st):

so good luck finding the zeros... and I hope there is a typo... I think there has to be 2 complex roots at a guess

OpenStudy (anonymous):

I'm so confused :/

OpenStudy (campbell_st):

is your original equation correct... \[x^3 - x^2 = -3x - 3\]

OpenStudy (anonymous):

Yes it is.

OpenStudy (anonymous):

I dunno. I have been stuck on 2 questions for about an hour now and they should be pretty easy but they aren't. :/

OpenStudy (anonymous):

\[i \sqrt{3} , 1\]

OpenStudy (anonymous):

Would that be the zeros?

OpenStudy (campbell_st):

nope... thats why I asked if the equation was correctly typed...

OpenStudy (anonymous):